OpenStudy (islandchick60):

Help! Will fan and medel! Which expression gives the solutions of -5+2x^2=-6x

OpenStudy (islandchick60):

These are the answer choices:

OpenStudy (islandchick60):

@likeabossssssss @MissSmartiez

OpenStudy (islandchick60):

@jomc4

OpenStudy (islandchick60):

@Compassionate

OpenStudy (islandchick60):

@dinamix ?

OpenStudy (dinamix):

\[(\Delta)' = b'^2-ac\]= 9+10=19 \[x=\frac{ -b'\pm \sqrt{\Delta'}}{ a } = \frac{ -4-\sqrt{19} }{ 2}\]

OpenStudy (islandchick60):

so... the third one?

OpenStudy (dinamix):

no

OpenStudy (islandchick60):

oh um i dont understand

OpenStudy (dinamix):

u know how to find root ?

OpenStudy (islandchick60):

u kinda

OpenStudy (dinamix):

so ?

OpenStudy (dinamix):

i mean rule to find it

OpenStudy (islandchick60):

what would i find the root for tho

OpenStudy (dinamix):

root of this equation

OpenStudy (islandchick60):

i am really confused

OpenStudy (dinamix):

 use this and u can find solution your self but just focus okey